# Optimizing reaction and transport fluxes in temperature gradient-driven chemical reaction-diffusion systems

### v1

@@ II.4 Coupling to a secondary reaction
Thus, energy can be directly extracted from a temperature gradient using these reaction-diffusion systems. However, the design of a coupled chemical reaction network implies the ability to propagate energy input to the secondary processes.[3, 14, 15] We further investigate the coupling of the minimalist network in Eq. (1) to maintain nonequilibrium secondary processes.
@@ III Conclusions
A general model was established to describe how an out-of-equilibrium steady reaction-diffusion process can be sustained by a temperature gradient. A difference in temperature across the system induces spatial differences in the chemical equilibrium states. This results in a frustrated state, because it is impossible to simultaneously sustain the chemical equilibrium (which would induce a concentration gradient) and the global transport equilibrium (which would imply homogeneous concentrations). Consequently, reaction-diffusion cycles are necessarily established, with chemical reactions counterbalancing chemical transports, as long as a temperature gradient is coupled to an endothermic or exothermic chemical process. We determined the parameters that enable the system to be dominated at will by either transport or reaction. Moreover, we demonstrated that the temperature gradient shape can be tuned such that both the reaction and transport processes can be either spatially disconnected from each other or concentrated in the same spatial area.
We showed that this nonequilibrium steady-state can be characterized in terms of:
  • chemical force, describing the local distance from equilibrium. This essentially originates from the reaction enthalpy, and the intensity of the temperature gradient.
@@
  • entropy production rate, describing the dissipative efficiency. This results from the combination of the chemical force and flux.
These characteristics can be understood as analogous to the electromotive force, current intensity, and dissipated power in an electric circuit, respectively. This yields a fruitful quantitative description of the nonequilibrium state. The expression of these characteristics was established analytically in the simple case of two homogeneous compartments that exchange matter . This result was then extended to interpret the complete reaction-diffusion systems, that were solved numerically.
The entropy dissipated by the reaction-diffusion system is distributed unequally among each process. A net difference can be observed between diffusion-dominated and reaction-dominated systems. The desired regime can be tailored by scaling the system to a critical system length $\mathcal{L}_{0}$. It is noteworthy that the traditional “local equilibrium” approximation reduces to the limit case of fast reactions, that is to large systems whose physical length is well above $\mathcal{L}_{0}$. This regime is suboptimal for generating nonequilibrium reaction fluxes; it is the worst case for extracting chemical energy from the temperature gradient, as it leads to minimization of the chemical flux.
@@
This study is based on an imposed steady temperature gradient. However, steady chemical reaction fluxes can act as secondary heat sources, that can influence back the temperature gradient from which they originates. Assuming a steady temperature gradient comes down to assume the presence of infinitely efficient external heat sources that can seamlessly absorb this chemical heat. Further work should describe the full coupling between the external and internal heat sources in these reaction-diffusion systems, to evaluate their impact. In addition, we should be able to explain how optimal reaction-mediated heat transport,[24] leading to an increase in thermal conductivity,[30] can be obtained.
@@ Acknowledgements.
This work was supported by the ANR grant "Sacerdotal" (ANR-19-CE6-0010-02). We thanks Alois Würger for the fruitful discussions on thermophoresis. We thank Nicolas Giuseppone, Nicolas Capit and Émilie Moulin for discussions on peptide exchange reactions.
@@
Additional information, as referred to in the main text is available in a supplementary PDF file. Simulation files (XMDS files, raw simulation data and processed data results) are available upon request.
@@ Appendix A Numerical tools
    A single unstirred system was connected to two spatially extended thermostats, each of which maintained a fixed temperature at each extremity over a length $\nicefrac{{1}}{{4}}$, leading to a steady linear gradient over the $[\nicefrac{{1}}{{4}},\nicefrac{{3}}{{4}}]$ central part of the system, with free chemical diffusion throughout the compartment.
The first model was solved by calculating the steady-state of the resulting ordinary differential equation (ODE) system. An analytical solution was obtained for a low gradient intensity. The full derivation of the steady-state behavior was performed using the Python symbolic computation package Sympy[42, 43]. It is available as a Jupyter notebook[44] in the SI, and the major results are detailed in Section C.
The other models were solved numerically by integrating the resulting partial differential equation (PDE) system (see section D.1). The steady-states of the reaction-diffusion systems were numerically computed for a wide set of values of $\beta\in[0.5,5]$, $\theta_{g}\in[0.01,0.5]$, $\rho\in[10^{-3},10^{+3}]$, $d\in[10^{-3},10^{+3}]$, $\gamma\in[0.2,1]$, $\kappa\in[0.1,10]$ and $\lambda\in[0.1,10]$. The steady-state was considered to be reached when the relative residual flux was sufficiently small, following:
@@
The computations were performed using XMDS2[45]. In each simulation, a steady temperature gradient was imposed with either a Heaviside, linear or mixed profile. In all models, the influence of temperature on the chemical reaction was introduced via the Arrhenius relationship for the reaction rate constants. Integration was performed using the 8th (embedded 9th) order adaptive Runge-Kutta method. Spatial derivation was performed based on a discrete cosine transform, for the implementation of zero Neumann boundaries, thus describing a closed system.
@@ Appendix B Parameter description
In non-stirred systems, the free diffusion of the chemical reactants is fully considered. Neglecting the variation in the diffusion constant with temperature, chemical diffusion is described as a function of the geometric mean of the diffusion constant $d$, and the exchange ratio $\lambda$:
@@ Appendix C Mathematical derivation in the stirred compartments model
C.1 steady-state
The steady-state $u_{ij}^{\infty}$ (with $i\in[1,2]$ and $j\in[a,b]$) can be directly obtained from the resolution of the set of equations (25aiaj), for $\dfrac{du_{ij}}{d\bar{t}}=0$. Each chemical force and flux of any reaction or exchange process $k$ can be expressed as:
@@
It is then possible to express the on-equilibrium steady-state of each reaction and compound exchange in terms of $[\alpha,\varphi]$ as:
@@
At the steady-state, all the chemical fluxes compensate for each other, following Eqs. (25aiaj) such that:
@@
Each term $\sigma_{k}$ is positive, so that $R_{k}\geq 0$; $\alpha_{k}$ and $\varphi_{k}$ are thus necessarily of the same sign.
@@
Solving the corresponding set of equations exactly leads to complex expressions of $\alpha_{i}$, $\varphi_{i}$ and $\sigma_{i}$ parameters. Thus, we simplified the system by assuming relatively small variations in temperature, using a second order Taylor expansion; $\alpha$ is further simplified to:
@@ C.2 Expression of the steady chemical flux
$\kappa^{\prime}$ and $\lambda^{\prime}$ are parameters that characterize the asymmetry of the chemical reaction and chemical exchange, respectively. The chemical flux in Eq. (25aiata) indicates that the system behavior depends on the relative values of $\rho$ and $\lambda^{\prime}\delta$, with the intensity proportional to the reaction asymmetry $\kappa^{\prime}$.
@@
Parameter $\varepsilon$ tends to $\beta$ when $\kappa\approx\lambda^{-1}$, to $\beta(1-\gamma)=\varepsilon_{+}$ when $\kappa$ or $\lambda$ tends to large values, and to $\beta(1+\gamma)=\varepsilon_{-}$ when $\kappa$ or $\lambda$ tends to zero. $\varepsilon$ can be interpreted as the global activation energy for the full system, varying from $\varepsilon_{+}$ to $\varepsilon_{-}$ through the average $\beta$ value, depending on the global symmetry of the system.
@@ C.5 General expression
In all situations, the steady-state can thus be described as:
@@
This results in a reference steady-state ($[\alpha_{0},\varphi_{0},\sigma_{0}]$), which depend on the thermokinetic characteristics of the chemical reaction. The state of a given system can be retrieved from this reference state based on Eqs. (25aibe).
@@ C.6 Entropy dissipation variations
For a given value of $\lambda^{\prime}\delta$, $\sigma_{r}$ is thus maximal for $\rho=\lambda\delta^{\prime}$; it varies proportionally to $\rho$ for $\rho\ll\lambda^{\prime}\delta$ and to $\rho^{-1}$ for $\rho\gg\lambda^{\prime}\delta$. For a given value of $\rho$, $\sigma_{r}$ monotonically increases with $\lambda^{\prime}\delta$; it varies proportionally to $(\lambda^{\prime}\delta)^{2}$ for $\lambda^{\prime}\delta\ll\rho$, and reaches a plateau for $\lambda^{\prime}\delta\gg\rho$.
@@
For a given value of $\rho$, $\sigma_{d}$ is thus maximal for $\lambda\delta^{\prime}=\rho$; it varies proportionally to $\lambda^{\prime}\delta$ for $\lambda^{\prime}\delta\ll\rho$ and to $(\lambda^{\prime}\delta)^{-1}$ for $\lambda^{\prime}\delta\gg\rho$. For a given value of $\lambda^{\prime}\delta$, $\sigma_{d}$ monotonically increases with $\rho$, varying proportionally to $\rho^{2}$ for $\rho\ll\lambda^{\prime}\delta$, but reaches a plateau for $\rho\gg\lambda^{\prime}\delta$ (see fig. S1 in SI.).
@@ D.1 Expression of the PDE model
In contrast to the two-compartments model, the variables are expressed as a function of the nondimensionalized spatial position $\bar{x}$, following an imposed temperature gradient $\theta(\bar{x})$, with $\theta(\bar{x})\in[-\theta_{g},+\theta_{g}])$.
The general properties of the system are described based on its characteristic concentration $c_{0}$ and temperature $T_{0}$. Howevere, the introduction of the spatial dimension introduces two new characteristics: length $l_{0}$ (corresponding to the system dimension), and diffusion constant $D_{0}$ (corresponding to the average diffusion constant of the reactants). The characteristic time is then directly obtained as $t_{0}=\nicefrac{{l_{0}^{2}}}{{D_{0}}}$. The system is thus described in terms of the nondimensionalized concentrations $u_{i}$, position $\bar{x}$, temperature deviation $\theta$, diffusion constants $d_{i}$ and time $\bar{t}$.
@@
D.2 Steady-state characteristics
The steady-state conditions must consider the spatial extension of the system. In the steady-state, the total concentrations of each compound $U_{1}$ and $U_{2}$ is constant. This implies that the global reaction is zero, i.e.:

### v2

@@ II.4 Coupling to a secondary reaction
Thus, energy can be directly extracted from a temperature gradient using these reaction-diffusion systems. However, the design of a coupled chemical reaction network implies the ability to propagate energy input to the secondary processes[3, 14, 15]. We further investigate the coupling of the minimalist network in Eq. (1) to maintain nonequilibrium secondary processes.
@@ III Conclusions
A general model was established to describe how an nonequilibrium steady reaction-diffusion process can be sustained by a temperature gradient. A difference in temperature across the system induces spatial differences in the chemical equilibrium states. This results in a frustrated state because it is impossible to simultaneously sustain the chemical equilibrium (which would induce a concentration gradient) and the global transport equilibrium (which would imply homogeneous concentrations). Consequently, reaction-diffusion cycles are necessarily established, with chemical reactions counterbalancing chemical transports, as long as a temperature gradient is coupled to an endothermic or exothermic chemical process. We determined the parameters that enable the system to be dominated at will by either transport or reaction. Moreover, we demonstrated that the temperature gradient shape can be tuned so that both the reaction and transport processes can be spatially disconnected from each other or concentrated in the same spatial area.
We showed that this nonequilibrium steady state can be characterized in terms of:
  • chemical force, describing the local distance from equilibrium. This essentially originates from the reaction enthalpy and the intensity of the temperature gradient.
@@
  • entropy production rate, which describes the dissipative efficiency. This results from the combination of chemical force and flux.
These characteristics can be understood as analogous to the electromotive force, current intensity, and dissipated power in an electric circuit, respectively. This yields a fruitful quantitative description of the nonequilibrium state. The expression of these characteristics was established analytically in the simple case of two homogeneous compartments that exchange matter. This result was then extended to interpret the complete reaction-diffusion systems that were solved numerically.
The entropy dissipated by the reaction-diffusion system is distributed unequally among each process. A net difference can be observed between the diffusion-dominated and reaction-dominated systems. The desired regime can be tailored by scaling the system to a critical system length $\mathcal{L}_{0}$. It should be noted that the traditional “local equilibrium” approximation reduces to the limit case of fast reactions, that is, to large systems whose physical length is well above $\mathcal{L}_{0}$. This regime is suboptimal for generating nonequilibrium reaction fluxes; it is the worst case for extracting chemical energy from the temperature gradient, as it leads to minimization of the chemical flux.
@@
This study is based on an imposed steady temperature gradient. However, steady chemical reaction fluxes can act as secondary heat sources that can influence back the temperature gradient from which they originate. Having a steady temperature gradient comes down to the assumption of the presence of infinitely efficient external heat sources that can seamlessly absorb this chemical heat. Further work should describe the full coupling between the external and internal heat sources in these reaction-diffusion systems, to evaluate their impact. In addition, we should be able to explain how optimal reaction-mediated heat transport[24], leading to an increase in thermal conductivity[30], can be obtained.
@@ Acknowledgements.
This work was supported by the ANR grant ”Sacerdotal” (ANR-19-CE6-0010-02). The authors thank Alois Würger for the fruitful discussions on thermophoresis. The authors thank Nicolas Giuseppone, Nicolas Capit, and Émilie Moulin for discussions on peptide exchange reactions.
@@
Additional information, as referred to in the main text, is available in a supplementary PDF file. Simulation files (XMDS files, raw simulation data, and processed data results) are available upon request.
@@ Appendix A Numerical tools
    A single unstirred system was connected to two spatially extended thermostats, each of which maintained a fixed temperature at each extremity over a length $\nicefrac{{1}}{{4}}$, leading to a steady linear gradient over the central part of the system $\bar{x}\in[\nicefrac{{1}}{{4}},\nicefrac{{3}}{{4}}]$, with free chemical diffusion throughout the compartment.
The first model was solved by calculating the steady state of the resulting ordinary differential equation (ODE) system. An analytical solution was obtained for a low gradient intensity. The full derivation of the steady state behavior was performed using the Python symbolic computation package Sympy[42, 43]. It is available as a Jupyter notebook[44] in the SI, and the major results are detailed in Section C.
The other models were solved numerically by integrating the resulting partial differential equation (PDE) system (see Section D.1). The steady states of the reaction-diffusion systems were numerically calculated for a wide set of values of $\beta\in[0.5,5]$, $\theta_{g}\in[0.01,0.5]$, $\rho\in[10^{-3},10^{+3}]$, $d\in[10^{-3},10^{+3}]$, $\gamma\in[0.2,1]$, $\kappa\in[0.1,10]$, and $\lambda\in[0.1,10]$. The steady state was considered to be reached when the relative residual flux was sufficiently small, following:
@@
The calculations were performed using XMDS2[45]. In each simulation, a steady temperature gradient was imposed with either a Heaviside, linear, or mixed profile. In all models, the influence of temperature on the chemical reaction was introduced via the Arrhenius relationship for the reaction rate constants. Integration was performed using the 8th (embedded 9th) order adaptive Runge-Kutta method. Spatial derivation was performed on the basis of a discrete cosine transform for the implementation of zero Neumann boundaries, thus describing a closed system.
@@ Appendix B Parameter description
In unstirred systems, the free diffusion of the chemical reactants is fully considered. Neglecting the variation in the diffusion constant with temperature, chemical diffusion is described as a function of the geometric mean of the diffusion constant $d$, and the exchange ratio $\lambda$:
@@ Appendix C Mathematical derivation in the stirred compartments model
C.1 steady state
The steady state $u_{ij}^{\infty}$ (with $i\in[1,2]$ and $j\in[a,b]$) can be obtained directly from the resolution of the set of equations (25aiaj), for $\dfrac{du_{ij}}{d\bar{t}}=0$. Each chemical force and flux of any reaction or exchange process $k$ can be expressed as:
@@
It is then possible to express the on-equilibrium steady state of each reaction and compound exchange in terms of $[\alpha,\varphi]$ as:
@@
At the steady state, all the chemical fluxes compensate for each other, following Eqs. (25aiaj) such that:
@@
Each term $\sigma_{k}$ is positive, so $R_{k}\geq 0$; $\alpha_{k}$ and $\varphi_{k}$ are thus necessarily of the same sign.
@@
Solving the corresponding set of equations exactly leads to complex expressions of $\alpha_{i}$, $\varphi_{i}$, and $\sigma_{i}$ parameters. Thus, we simplified the system by assuming relatively small variations in temperature, using a second order Taylor expansion; $\alpha$ is further simplified to:
@@ C.2 Expression of the steady chemical flux
$\kappa^{\prime}$ and $\lambda^{\prime}$ are parameters that characterize the asymmetry of the chemical reaction and chemical exchange, respectively. The chemical flux in Eq. (25aiata) indicates that the behavior of the system depends on the relative values of $\rho$ and $\lambda^{\prime}\delta$, with the intensity proportional to the reaction asymmetry $\kappa^{\prime}$.
@@
The parameter $\varepsilon$ tends to $\beta$ when $\kappa\approx\lambda^{-1}$, to $\beta(1-\gamma)=\varepsilon_{+}$ when $\kappa$ or $\lambda$ tends to large values and to $\beta(1+\gamma)=\varepsilon_{-}$ when $\kappa$ or $\lambda$ tends to zero. $\varepsilon$ can be interpreted as the global activation energy for the whole system, varying from $\varepsilon_{+}$ to $\varepsilon_{-}$ through the average value of $\beta$, depending on the global symmetry of the system.
@@ C.5 General expression
In all situations, the steady state can thus be described as:
@@
This results in a reference steady state ($[\alpha_{0},\varphi_{0},\sigma_{0}]$), which depends on the thermokinetic characteristics of the chemical reaction. The state of a given system can be retrieved from this reference state based on Eqs. (25aibe).
@@ C.6 Entropy dissipation variations
For a given value of $\lambda^{\prime}\delta$, $\sigma_{r}$ is thus maximal for $\rho=\lambda\delta^{\prime}$; it varies proportionally to $\rho$ for $\rho\ll\lambda^{\prime}\delta$ and to $\rho^{-1}$ for $\rho\gg\lambda^{\prime}\delta$. For a given value of $\rho$, $\sigma_{r}$ increases monotonically with $\lambda^{\prime}\delta$; it varies proportionally to $(\lambda^{\prime}\delta)^{2}$ for $\lambda^{\prime}\delta\ll\rho$, and reaches a plateau for $\lambda^{\prime}\delta\gg\rho$.
@@
For a given value of $\rho$, $\sigma_{d}$ is thus maximal for $\lambda\delta^{\prime}=\rho$; it varies proportionally to $\lambda^{\prime}\delta$ for $\lambda^{\prime}\delta\ll\rho$ and to $(\lambda^{\prime}\delta)^{-1}$ for $\lambda^{\prime}\delta\gg\rho$. For a given value of $\lambda^{\prime}\delta$, $\sigma_{d}$ increases monotonically with $\rho$, varying proportionally to $\rho^{2}$ for $\rho\ll\lambda^{\prime}\delta$, but reaches a plateau for $\rho\gg\lambda^{\prime}\delta$ (see fig. S1 in SI.).
@@ D.1 Expression of the PDE model
In contrast to the two-compartment model, the variables are expressed as a function of the nondimensionalized spatial position $\bar{x}$, following an imposed temperature gradient $\theta(\bar{x})$, with $\theta(\bar{x})\in[-\theta_{g},+\theta_{g}])$.
The general properties of the system are described based on its characteristic concentration $c_{0}$ and temperature $T_{0}$. However, the introduction of the spatial dimension introduces two new characteristics: length $l_{0}$ (corresponding to the system dimension), and diffusion constant $D_{0}$ (corresponding to the average diffusion constant of the reactants). The characteristic time is then directly obtained as $t_{0}=\nicefrac{{l_{0}^{2}}}{{D_{0}}}$. The system is thus described in terms of the nondimensionalized concentrations $u_{i}$, position $\bar{x}$, temperature deviation $\theta$, diffusion constants $d_{i}$, and time $\bar{t}$.
@@
D.2 steady state characteristics
The steady-state conditions must consider the spatial extension of the system. In steady state, the total concentrations of each compound $U_{1}$ and $U_{2}$ are constant. This implies that the global reaction is zero, i.e.: